what is the solution of the equation e^xe^2x=4?
Answer by stanbon(75887)

what is the solution of the equation e^xe^2x=4?
---------------------------
e^(3x) = 4
Take the natural log to get:
3x = ln(4)
x = (1/3)ln(4)
x = 0.4621...
Cheers,
Stan H.